• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name folts

    Log entry time 16:17:11 on June 25, 2010

    Entry number 328190


    The checklist is complete

    Tech on call at startup is Todd Ewing

    Techs are to be called for emergencies only

    Left sieve position is out

    Right sieve position is out

    Left collimator position is out

    Right collimator position is out

    Septum magnet is operational

    Moeller is not operational

    Target chamber vacuum is <7X10-5 and dropping

    The right detector stack is out

    The right spectrometer is not operational

    It needs to be noted that several red flashing beacons have burned out due to the radiation including the one over the septum magnet.

    There is a lot of ice around the leads of the right Q1 magnet and some of the heaters do not seem to be working.